By loosening its credit standards, the CFO of the Henleigh Company believes she can increase the net present value of the firm’s daily cash flows by $10,000. If the increased daily NPV continues indefinitely, what is the aggregate net present value of the decision to lessen the credit standards? Assume a 365-day year. Henry uses a 8% cost of capital for credit policy decisions.
Answer #1

IT IS A CASE OF PERPETUITY.

PRESENT VALUE = ANNUAL AMOUNT/ RATE

BUT AS AMOUNT IS GIVEN IN DAILY TERMS, WE HAVE TO CONVERT IN DAILY TERMS

PRESENT VALUE = DAILY AMOUNT/DAILY RATE

DAILY AMOUNT = 10000, DAILY RATE = 8%/365 = 0.08/365

PRESENT VALUE = DAILY AMOUNT/DAILY RATE = 10000/ (0.08/365) = 45,625,000

ANSWER : $45,625,000 (Thumbs up please)